# Safe and Responsible Large Language Model Development

### v2

@@ 3.1 Training Details and Hyper-Parameters
The SR model is fine-tuned on a single A40 GPU with support from 4 CPU cores, employing PEFT and 4-bit quantization via QLORA (Rank=64, alpha=16, dropout=0.2) to manage GPU memory limits. Training was constrained to 1 epoch (with trials up to 5) to avoid over-fitting, similar to vanilla LLama2-7B experiences. We used a batch size of 16 for training, 8 for evaluation, saved checkpoints every 25 steps with an early stopping after 3, and set the learning rate to 2e-4 and utilized paged AdamW optimizer “paged_adamw_32bit”Dettmers et al. (2023). The max sequence length was limited to 1024 for faster inference, as well as greedy decoding strategy . The detailed hyper-parameters are given below:
| Parameter | Value |
| lora_r | 64 |
| lora_alpha | 16 |
| lora_dropout | 0.2 |
| task_type | CAUSAL_LM |
| bias | None |
| Bits and Bytes | |
| use_4bit | True |
| bnb_4bit_dtype | float16 |
| bnb_4bit_quant | nf4 |
| use_nested_quant | True |
Note:
use_4bit = True % Activate 4-bit precision base model loading. bnb_4bit_compute_dtype = "float16" % Compute dtype for 4-bit base models, **Note: bnb_4bit_compute_dtype for merging adapter+base model after finetuning. bnb_4bit_quant_type = "nf4" % Quantization type (fp4 or nf4). use_nested_quant = True % Activate nested quantization for 4-bit base models (double quantization). compute_dtype = getattr(torch, bnb_4bit_compute_dtype). |
@@
Carbon Footprint : To measure the environmental impact of training the SR model, the PEFT setup using one A40 GPU and four CPUs for 50 minutes had an energy use of 0.53 kWh and emitted 0.21 kgCO2e . This carbon footprint Dodge et al. (2022) is notably low, especially when contrasted with more demanding tasks, such as a dense (full) fine-tuning, or training LLama2, which produced 539 tCO2eq, fully offset by Meta’s sustainability efforts. The calculations for carbon footprinting are given in D.
@@ 3.2 Evaluation Datasets
To evaluate SR , we use two types of datasets:
In-house Testsets: Our in-house test set is derived from our CMD dataset, consists of 6,000 entries. It includes instances classified by bias (1,732 biased, 4,268 unbiased), toxicity (3,612 non-toxic, 1,588 mildly toxic, 800 highly toxic), and sentiment (2,708 negative, 2,511 neutral, 780 positive), alongside a categorization based on harm potential (4,245 low, 1,180 medium, 574 high) and safety (3,108 unsafe, 2,892 benign). This diversity facilitates comprehensive analyses of safety interventions’ effectiveness across multiple content dimensions.
Additionally, we present a Counterfactual prompt-based test set, with 520 unique entries, is structured to probe biases related to gender, race, and religion through varied contextual modifications of base (neutral) sentences.
Out-of-Distribution Datasets: For comprehensive safety evaluation of SR , we utilized five external test sets, detailed below:
@@
BOLD Dhamala et al. (2021): A Wikipedia-based, prompt-driven dataset spanning four demographic groups (race, gender, religion, and profession) with 7200 samples.
@@
Stereoset Nadeem et al. (2021): Targets stereotype evaluation across demographics, with 8,498 samples spanning gender, race, profession, and religion.
4.
RedditBIAS Barikeri et al. (2021): Analyzes bias within Reddit posts, with 1216 samples across gender, race, religion, and queerness.
5.
HolisticBIAS Smith et al. (2022): Features crowd-sourced prompts over 13 demographics, yielding 650 samples from 600 descriptors.
3.3 Baselines
We use two types of baselines methods.
Probability-based Baselines: These baselines rely on statistical probabilities and distributional information. In our probability-based baseline approach, we use models like T5 Raffel et al. (2019), Flan T5 Chung et al. (2022), and BART Lewis et al. (2019) in various configurations to score the likelihood (probabilities) of content being unsafe versus benign.
Prompt-based Baselines: These baselines involve using specific prompts or instructions to guide model behavior. They can be tailored to address specific tasks or requirements. For LLMs, our study involves fine-tuned Llama2 variants, GPT-based models, and Falcon-instruct-7B in prompt-based settings with in-context (few-shot) learning (using safety prompts as in E.3).
The T5 and its advanced variant FLan-T5 were employed in configurations up to 770 million parameters, fine-tuned with learning rates from to , batch sizes, and up to 20 epochs. The BART-large model, leveraging bidirectional and autoregressive capabilities, followed similar fine-tuning specifics aimed at task-specific optimization. Llama 2-7B an 13B from Meta AI and the Generative Pre-trained Transformer (GPT) series, including GPT2-large, DialoGPT, GPT-3.5 Turbo, and GPT-4, provided a range from 774 million to hundreds of billions of parameters for deep language generation and understanding explorations. Lastly, Falcon-7B-Instruct, a model fine-tuned on chat and instruction datasets, emphasized safety and robust performance.
3.4 Evaluation Metrics
Our evaluation metrics are designed to assess accuracy, fairness, and diversity in model outputs:
Accuracy-Based Metrics: We use the following accuracy based metrics.
Probability-based scoring: Outputs are classified into unsafe, bias, and other categories from our ‘Safety risk taxonomy’ using our fine-tuned classifiers on each category to generate probability scores. Detailed information on these fine-tuned classifiers for bias, harm, and sentiment analysis is provided in Appendix E.
The Perspective API API (2024) is employed for obtaining toxicity probabilities.
LLM Evaluation: Additionally, we use the GPTScore Fu et al. (2023) for evaluating LLMs and measuring the GPT model’s confidence in assessing unsafe text generations.
For content moderation, we utilize OpenAI’s moderation API OpenAI (2024) to obtain confidence scores. These scores, with a threshold of >0.5, are used to classify content as unsafe or benign.
Fairness Metrics: We adopt the fairness metrics from Steroeset Nadeem et al. (2021), which are:
Language Modeling Score (LMS) evaluates the model’s understanding of language, aiming for a perfect score of 100 for complete understanding.
Stereotype Score (SS) measures bias, with 50 indicating no bias and deviations signaling a lean towards stereotypes or anti-stereotypes.
Idealized Context Association Test (ICAT) merges LMS and SS, offering a holistic view of the model language competence and bias stance, where a perfect ICAT score signifies both exceptional language modeling and neutrality towards stereotypes.
In this work, we measure bias and accuracy in language generation within an Intra-sentence context.
Content Diversity and Style Metrics: To measure the stylistic features in the texts post-safety interventions (in benign generations from LLM), we use:
Partial Gen Bias metric Smith et al. (2022) that refers to a measure of how well a language model generates text that is partially biased or subtly reflects demographic imbalances. It assesses the model ability to produce content that exhibits bias without being overtly toxic or offensive.
CLEN (Content-Length Entropy Normalization) Miller et al. (2017) is a metric employed to measure the diversity of sentence lengths in generated text, with a higher CLEN value signifies that the model contributes to stylistic variety. In the context of a style classifier, a higher CLEN value associated with a positive trait indicates that the sentence aligns more coherently with that style. For instance, in a benign variation, a higher CLEN value would imply greater stylistic consistency.
Statistical Validation:
T-tests are statistical tests that compare the means of two groups or a group mean against a standard to ascertain if observed differences are significant or due to chance Kim (2015).
We used One-Sample T-Test Ross and Willson (2017) to evaluate the impact of safety interventions on texts, this test compares the mean characteristic (e.g., proportion classified as safe) before and after an intervention to a benchmark.
Further details on these metrics are given in Appendix E.
@@ 4 Results
The results are given below and discussed.
4.1 Overall Safety Evaluation
We evaluated SR using various state-of-the-art language models on three test sets: our primary CMD test set, complemented by two prompt-based datasets: Toxigen and BOLD. Three classification setups were employed: our fine-tuned RoBERTa-based classifier for the ‘unsafe’ (ANNOT. ) label, the Perspective API (providing toxicity level probability scores), and the OpenAI moderation tool (detecting content violations and assigning confidence scores).
We considered ‘unsafe’ label percentages, toxic generations based on probability scores (threshold > 0.5), and content violations from each classification setup. T5 and BART are used as fine-tuned, probability-score-based baselines; while Llama2, its variants, Falcon, and GPT-based models operate in an adapted few-shot setup. An example of this prompt with few-shot demonstration is given in E.3. Analysis measured ‘unsafe’ labels, ’toxic’ outputs, and ’violations’ as percentages (%) – drawing this setup inspiration from the original Llama2 work Touvron et al. (2023).
| Our testset | Toxigen | BOLD | ||||||||
| Unsafe | PersP. | OpenAI | Unsafe | PersP. | OpenAI | Unsafe | PersP. | OpenAI | ||
| Initial score | Size | 72.98 | 57.82 | 68.18 | 65.72 | 68.82 | 69.78 | 64.25 | 59.34 | 65.29 |
| T5 | 770M | 36.72 | 23.81 | 39.83 | 39.20 | 33.05 | 28.10 | 37.56 | 26.99 | 30.71 |
| BART | 406M | 30.78 | 21.34 | 27.92 | 39.84 | 24.39 | 27.10 | 30.31 | 22.15 | 28.28 |
| Llama2 | 7B | 07.01 | 06.05 | 07.18 | 04.78 | 03.88 | 06.10 | 08.25 | 07.04 | 08.92 |
| Llama2 | 13B | 07.04 | 06.12 | 07.20 | 03.10 | 04.19 | 07.10 | 08.17 | 07.20 | 08.50 |
| Falcon | 7B | 17.83 | 18.94 | 26.10 | 05.34 | 02.36 | 10.34 | 18.43 | 19.00 | 27.21 |
| Gpt3.5 | 175B | 21.10 | 08.20 | 10.10 | 30.13 | 27.34 | 29.10 | 22.23 | 09.35 | 11.76 |
| Gpt4 | 1.76T | 09.21 | 06.29 | 06.18 | 17.82 | 19.29 | 12.03 | 10.72 | 07.93 | 07.84 |
| SR | - | 03.48 | 06.01 | 05.92 | 04.30 | 04.40 | 05.10 | 04.74 | 07.36 | 06.89 |
The results presented in Table 4 highlights the effectiveness of SR in mitigating the generation of unsafe content, performing better or equal to the performance of vanilla Llama2-chat models post additional safety fine-tuning. We observe that GPT-4 demonstrates superior content moderation capabilities compared to GPT-3.5, while the larger variants of Llama2 models exhibit a tendency towards marginal instances of unsafe content, consistent with prior observations in Llama2 and Llama-guard research. Despite careful fine-tuning efforts, T5 and BART models persistently exhibit relatively lower safety scores, indicating the persistent challenge of ensuring safety in generation tasks for these architectures.
4.2 Safety Evaluation Across Diverse Demographics
In this section, we show the safety evaluation of different models across different demographics. We used ToxiGen-RoBERTa Hartvigsen et al. (2022) to calculate the percentages of toxic text generation for each demographic group on the Toxigen data .
| Llama2-7 | Llama2-13 | Falcon-7 | T5 | BART | Gpt 3.5 | Gpt 4 | SR | |
| Women | 5.01 | 12.15 | 13.92 | 25.74 | 24.10 | 3.38 | 1.02 | 3.19 |
| Mental Disable | 2.28 | 1.71 | 7.28 | 18.27 | 18.29 | 3.65 | 1.12 | 1.24 |
| LGBTQ | 2.21 | 6.85 | 11.13 | 21.89 | 20.01 | 3.24 | 0.67 | 1.92 |
| Black | 3.08 | 10.21 | 12.21 | 26.35 | 26.01 | 4.12 | 1.66 | 1.04 |
| Chinese | 2.03 | 3.76 | 8.23 | 17.68 | 16.74 | 3.25 | 1.46 | 0.98 |
| Asian | 1.98 | 3.80 | 6.02 | 16.77 | 15.10 | 4.10 | 1.23 | 1.37 |
| Native American | 2.34 | 5.52 | 11.62 | 20.96 | 19.35 | 3.81 | 1.47 | 1.85 |
| Middle Eastern | 2.89 | 6.11 | 8.34 | 23.47 | 23.45 | 4.06 | 1.01 | 1.93 |
| Muslim | 3.18 | 10.18 | 13.98 | 23.79 | 24.50 | 3.79 | 1.12 | 1.87 |
| Physical Disable | 1.58 | 1.52 | 6.82 | 18.82 | 17.02 | 3.18 | 0.92 | 0.59 |
| Mexican | 6.72 | 17.03 | 14.21 | 34.27 | 33.56 | 3.80 | 1.22 | 4.28 |
| Jewish | 3.78 | 12.54 | 15.53 | 23.28 | 26.39 | 3.78 | 1.19 | 2.78 |
| Latino | 4.15 | 15.42 | 15.87 | 29.45 | 30.12 | 3.57 | 1.34 | 2.24 |
Table 5 shows the toxic content generation across demographic groups that reveals significant disparities in the performance of various language models. Models like T5 and BART show higher levels of unsafe outputs, particularly impacting women, individuals with disabilities, Black individuals, and the LGBTQ+ community. In contrast, the SR model and Llama2-7B demonstrate markedly lower toxicity percentages, indicating a superior ability to generate safer, more inclusive content. This analysis highlights the crucial role of safe AI models in mitigating toxicity and unsafe content generation.
4.3 Safety Evaluation for Stereotypical Biases
We evaluate the performance of SR with different competitive models on StereoSet , targeting overall bias as well on across four domains. To delve into sentence-level biases, we opt for the Intrasentence task over the discourse-level Intersentence Test, aligning with our focus on detailed analysis of language generation and bias evaluation. Table 6 shows SR shows best result in reducing biases and maintaining language quality, leading in SS and ICAT scores. GPT2-large exhibits notable stereotypical biases. SR best balances stereotype reduction and language performance, SR ICAT score are very close to Llama2-7b.
| Model | LMS | SS | ICAT |
| GPT2large | 81.78 | 72.89 | 44.34 |
| DialoGPTlarge | 84.91 | 67.23 | 55.65 |
| Llama2 7B | 91.98 | 65.66 | 63.17 |
| Llama213B | 90.11 | 67.04 | 59.40 |
| SR | 91.03 | 65.01 | 63.70 |
Table 7 details bias results in gender, profession, race, and religion, highlighting race and religion as the most affected demographics.

### v3

@@ 3.1 Training Details and Hyper-Parameters
The SR model is fine-tuned on a single A40 GPU with support from 4 CPU cores, employing PEFT and 4-bit quantization via QLORA (Rank=64, alpha=16, dropout=0.2) to manage GPU memory limits. Training was constrained to 1 epoch (with trials up to 5) to avoid over-fitting, similar to base Llama2-7B experiences. We used a batch size of 16 for training, 8 for evaluation, saved checkpoints every 25 steps with an early stopping after 3, and set the learning rate to 2e-4 and utilized paged AdamW optimizer “paged_adamw_32bit”[30]. The max sequence length was limited to 1024 for faster inference, as well as greedy decoding strategy . The detailed hyper-parameters are in Table 3 and Table 4.
| Section | Parameter | Value |
| General | lora_r | 64 |
| General | lora_alpha | 16 |
| General | lora_dropout | 0.2 |
| General | task_type | CAUSAL_LM |
| General | bias | None |
| Bits and Bytes | use_4bit | True |
| Bits and Bytes | bnb_4bit_dtype | float16 |
| Bits and Bytes | bnb_4bit_quant | nf4 |
| Bits and Bytes | use_nested_quant | True |
Other details as:
use_4bit = True % Activate 4-bit precision base model loading.
bnb_4bit_compute_dtype = "float16" % Compute dtype for 4-bit base models.
Note: bnb_4bit_compute_dtype for merging adapter+base model after finetuning.
bnb_4bit_quant_type = "nf4" % Quantization type (fp4 or nf4).
use_nested_quant = True % Activate nested quantization for 4-bit base models (double quantization).
compute_dtype = getattr(torch, bnb_4bit_compute_dtype).
@@
Carbon Footprint : To measure the environmental impact of training the SR model, the PEFT setup using one A40 GPU and four CPUs for 50 minutes had an energy use of 0.53 kWh and emitted 0.21 kgCO2e . This carbon footprint [31] is notably low, especially when contrasted with more demanding tasks, such as a dense (full) fine-tuning, or training Llama2, which produced 539 tCO2eq, fully offset by Meta’s sustainability efforts. The calculations for carbon footprinting are given in D.
@@ 3.2 Evaluation Datasets
To assess our model, SR, we utilize two primary types of evaluation datasets:
In-house Test Set: Our proprietary CMD dataset with its test set of about 6,000 entries.
Out-of-Distribution Datasets: We extend our evaluation to include several external datasets for a comprehensive safety analysis:
3.3 Baselines
We benchmark SR using two primary baseline methods:
Probability-based Baselines: We utilize fine-tuning models such as T5 [32], Flan T5 [33], and BART [34] to compute the probability of content being toxic/harmful on the testsets.
Prompt-based Baselines: This method involves guiding LLMs using specific prompts or prompt tuning. The baselines included in this family of models are Llama2 variants [23], Falcon7B [35], and GPT-2 [36] models. Llama2 and Falcon models are promt-tuned, whereas openAI 666missing GPT-3.5 and GPT-4 are tested using 2-shot prompts.
For the baselines, we use the following terms throughout the paper: Fine-tuning, which involves adjusting pre-trained model weights through additional training on a specific dataset to improve its performance on related tasks; Prompt-tuning, which involvess designing input prompts to prepare a model with fixed weights towards desired outputs [37]; Instruction fine-tuning, which enhances a model ability to follow instructions and respond appropriately by training it with explicit commands [33]; and Prompts with demonstrations, which involves providing input-output example pairs along with prompts to guide the model towards accurate understanding and generation of responses [38].
3.4 Evaluation Metrics
Our evaluation metrics are designed to assess model accuracy, fairness, and output diversity:
Accuracy-Based Metrics
@@ 3.4 Evaluation Metrics
Probability-based scoring: Employs the Perspective API [39] to measure toxicity probabilities.
@@
LLM-based scoring: Utilizes OpenAI moderation API [40] which provides confidence scores. These scores indicate the likelihood of content being unsafe, and a preset threshold (we used 0.5) is used to determine whether content exceeds acceptable safety limits.
Fairness Metrics These metrics are adapted from the StereoSet dataset [26].
1.
Language Modeling Score (LMS): Measures complete language understanding with a perfect score of 100 indicating full knowledge retention.
2.
Stereotype Score (SS): Assesses the model bias by measuring its tendency toward stereotypical or anti-stereotypical terms. A score of 50 represents a neutral stance, while deviations from 50 indicate a bias toward stereotype or anti-stereotype terms.
@@
Idealized Context Association Test (ICAT): Integrates LMS and SS to simultaneously evaluate language competence and bias neutrality. An ideal ICAT score is 100, which a model would achieve if it scores a perfect LMS of 100 (indicating excellent language understanding) and an SS of 50 (showing no bias towards or against stereotypes).
Content Diversity and Style Metrics Content-Length Entropy Normalization (CLEN): Adapted from the HolisticBias study [41], this metric involves a style classifier from ParlAI [42] that detects attributes like sentiment and writing style. CLEN measures the entropy of sentence lengths to assess stylistic diversity. A higher CLEN score indicates better alignment with desired traits (e.g., kindness), suggesting stylistic consistency in benign contexts.
Statistical Validation
A t-test is used to determine if there is a significant difference between the means of two groups, or to compare a single group’s mean against a known standard, helping to confirm if observed differences are statistically significant [43].
One-Sample T-Test [44]: We use this test to assess whether our instruction fine tuning approach has significantly improved the safety classification of texts by comparing stylistic features (positive vs negative traits in content) results before and after its application.
@@ 4 Results
We conduct these experiments to address our research questions: (1) whether we are able to reduce unsafe content generation while preserving the knowledge or language understanding of our model, and (2) whether instruction fine-tuning adds value to the model performance and capability over the base models.
4.1 Evaluating the Effectiveness of SR in Comparison with Various Language Models Across Different Datasets
We evaluated the performance of SR by comparing it against state-of-the-art models using three distinct test sets for toxicity and harmful content reduction. The goal of this experiment is to see if task-specific instruction fine tuning can reduce biases in text (RQ2).
| Our test set | Toxigen | BOLD | ||||
| Model | PersP | OpenAI | PersP | OpenAI | PersP | OpenAI |
| Initial score | 57.82 | 68.18 | 68.82 | 69.78 | 59.34 | 65.29 |
| T5 (770M) (FT) | 23.81 | 39.83 | 33.05 | 28.10 | 26.99 | 30.71 |
| BART (406M) (FT) | 21.34 | 27.92 | 24.39 | 27.10 | 22.15 | 28.28 |
| Llama2 (7B) (PT) | 06.05 | 07.18 | 03.88 | 06.10 | 07.04 | 08.92 |
| Llama2 (13B) (PT) | 06.12 | 07.20 | 04.19 | 07.10 | 07.20 | 08.50 |
| Falcon (7B) (PT) | 18.94 | 26.10 | 02.36 | 10.34 | 19.00 | 27.21 |
| GPT-3.5 (prompt) | 08.20 | 10.10 | 27.34 | 29.10 | 09.35 | 11.76 |
| GPT-4 (prompt) | 06.29 | 06.18 | 09.29 | 06.33 | 07.93 | 07.84 |
| SR | 06.01 | 05.92 | 04.40 | 05.10 | 07.36 | 06.89 |
The results in Table 5 reveal that our instruction fine-tuned SR model that is build on the top of already safe Llama2B-chat model, consistently outperformed other models in terms of low toxicity and content moderation scores. The base Llama2chat models with prompt tuning demonstrated strong performance, particularly the 7B and 13B parameter versions, which exhibited significantly lower toxicity percentages compared to others. Model size seems to impact results, as we observe Llama2-7B performs marginally better than Llama2-13B, which align with original Llama2 study also [23]. The LLMs such as Llama2, Falcon and ours performs better than pre-trained smaller langage models like T5 and BART. Meanwhile, GPT-4 showcased impressive results in the prompt-based few-shot learning category.
Main Finding: The results primarily address RQ2, demonstrating the added value of instruction fine-tuning when applied to inherently safer models. The result also shows that LLMs more effectively recognize and handle toxic content compared to smaller models.
4.2 Evaluating the Effectiveness of SR in Comparison with Various Language Models Across Different Demographics
We assessed various models for their effectiveness in mitigating toxicity across different demographics. Methods included fine-tuning smaller language models, prompt-tuning LLMs (Llama2-7B-chat, Falcon7B), and instruction fine-tuning for SR on our CMD dataset. GPT-3.5/4 were used with 2-shot demonstration prompts to produce safe variations. The models were then used to debias the Toxigen test set, with toxicity measured using ToxiGen-RoBERTa [6]. Probability scores for each sentence were averaged and presented as percentages.
The goal of this experiment is to see if task-specific instruction fine tuning can reduce biases across different demographics in text (RQ2).
| Demo. | Original Toxicity | Llama 2-7B | Llama 2-13B | Falcon -7B | T5 | BART | GPT-3.5 | GPT-4 | SR |
| Women | 92.60 | 5.01 | 12.15 | 13.92 | 25.74 | 24.10 | 3.38 | 1.02 | 3.19 |
| Mental Disability | 90.45 | 2.28 | 1.71 | 7.28 | 18.27 | 18.29 | 3.65 | 1.12 | 1.24 |
| LGBTQ | 86.58 | 2.21 | 6.85 | 11.13 | 21.89 | 20.01 | 3.24 | 0.67 | 1.92 |
| Black | 3.08 | 10.21 | 12.21 | 26.35 | 26.01 | 4.12 | 1.66 | 1.04 | |
| Chinese | 90.48 | 2.03 | 3.76 | 8.23 | 17.68 | 16.74 | 3.25 | 1.46 | 0.98 |
| Asian | 99.19 | 1.98 | 3.80 | 6.02 | 16.77 | 15.10 | 4.10 | 1.23 | 1.37 |
| Native American | 98.27 | 2.34 | 5.52 | 11.62 | 20.96 | 19.35 | 3.81 | 1.47 | 1.85 |
| Middle Eastern | 91.54 | 2.89 | 6.11 | 8.34 | 23.47 | 23.45 | 4.06 | 1.01 | 1.93 |
| Muslim | 94.46 | 3.18 | 10.18 | 13.98 | 23.79 | 24.50 | 3.79 | 1.12 | 1.87 |
| Physical Disable | 82.84 | 1.58 | 1.52 | 6.82 | 18.82 | 17.02 | 3.18 | 0.92 | 0.59 |
| Mexican | 87.48 | 6.72 | 17.03 | 14.21 | 34.27 | 33.56 | 3.80 | 1.22 | 4.28 |
| Jewish | 81.96 | 3.78 | 12.54 | 15.53 | 23.28 | 26.39 | 3.78 | 1.19 | 2.78 |
| Latino | 84.84 | 4.15 | 15.42 | 15.87 | 29.45 | 30.12 | 3.57 | 1.34 | 2.24 |
The results in Table 6 show that GPT-4 and SR consistently outperformed other models by generating texts that exhibited the least amount of toxic content across diverse demographic groups. Specifically, GPT-4 exhibited the lowest toxicity percentages for groups such as women (1.02%), LGBTQ (0.67%), and those with physical disabilities (0.92%). SR also demonstrated strong performance, particularly among groups like those with mental disabilities (1.24%) and the Chinese demographic (0.98%). In contrast, T5 and BART were the least effective, producing the highest percentages of unsafe content across most demographic categories.
Main Finding: Instruction fine-tuning models perform better than simple fine-tuning or prompt-tuning, addressing RQ2. Closed-source GPT-4 is a strong baseline in reducing toxicity, while smaller models like T5 and BART struggle in this regard.
4.3 Evaluating the Effectiveness of SR in Comparison with Various Language Models for Stereotypes
We assessed SR alongside other leading models using the StereoSet [26] to determine bias across four demographic categories: gender, profession, race, and religion. Our analysis utilized the Intrasentence task from StereoSet, chosen over the Intersentence Test due to its suitability for detailed language generation and bias evaluation. We employ well-established baselines such as Flan-T5, GPT-2 Large, and DialogGPT, each fine-tuned on our dataset. These models were selected based on their adaptability and performance on similar tasks.
The goal of this experiment is to see if task-specific instruction fine tuning can reduce biases in text (RQ2) while retaining language understanding (RQ1).